Let sin m = .11. Which of the following is sin m/2?
1
Expert's answer
2013-07-24T07:31:49-0400
There's formula relating sin m/2 and cos m:
sin m/2=sqrt((1-cos m)/2)
Also cos m=sqrt(1-sin^2 m)=sqrt(1-0.11^2)=0.99
So sin m/2=sqrt((1-cos m)/2)=sqrt((1-0.99)/2)=0.07